# Foldings in relatively hyperbolic groups

### v1

@@ Introduction
Makanin-Razborov diagrams refine the above arguments as they give a complete parametrization of all homomorphisms from a given finitely presented (or even finitely generated) group to a given (toral relatively) hyperbolic group. They rely essentially on the ideas of Makanin [34] and Razborov [37] and are in their modern form due to Sela [40] and Kharlampovich-Myasnikov [31, 32] in the case of free groups, see [41] and [38] for hyperbolic groups and [25, 24] for toral relatively hyperbolic groups.
@@
Suppose that $G$ is a group that is hyperbolic relatively to a collection of subgroups $\{P_{1},\ldots,P_{n}\}$. The following three settings can (and often do) yield infinitely many conjugacy classes of one-ended isomorphic subgroups of $G$:
@@
    Let $H=H_{1}*_{C}H_{2}$ be a subgroup of $G$ such that $C\leq P_{i}$ for some $i$ and let $p\in P_{i}\setminus C$ such that $p$ centralizes $C$. Then $\varphi_{p}:H\to G$ given by $\varphi_{p}(h_{1})=h_{1}$ for all $h_{1}\in H_{1}$ and $\varphi_{p}(h_{2})=ph_{2}p^{-1}$ for all $h_{2}\in H_{2}$ defines a homomorphism. It can often be ensured using a combination theorem of Martinez-Perdrosa that $\varphi_{p}$ is injective and that there are infinitely many such $p$ such that images of the $\varphi_{p}$ yield isomorphic and pairwise non-conjugate subgroups isomorphic to $H$.
@@
    Let $H=H_{1}*C=\langle H,t\mid tct^{-1}=\psi(c)\hbox{ for all }c\in C\rangle$ where $C\subset H_{1}\subset G$ and $C\subset P_{i}$. Let $p\in P_{i}\setminus C$ such that $p$ centralizes $C$. Then $\varphi_{p}:H\to G$ given by $\varphi_{p}(h_{1})=h_{1}$ for all $h_{1}\in H_{1}$ and $\varphi_{p}(t)=tp$ defines a homomorphism. As in (1) this construction often yields infintely many isomorphic and pairwise non-conjugate subgroups isomorphic to $H$.
@@
This is achieved by the following three steps, here $(G,\mathbb{P})$ is a torsion-free group $G$ that is hyperbolic relative to a collection of subgroups =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$.
@@
In chapter one we recall basic notions for relatively hyperbolic groups and establish a sufficient condition for a word to be a quasigeodesic. In Chapter 2 we introduce $(G,\mathbb{P})$ carrier graphs of groups and study their basic properties before we prove the combination theorem in Section 3. Section 4 is then dedicated to the introduction of folds and the proofs of the finiteness theorems. We conclude the paper with a brief discussion of the case of relatively hyperbolic groups with torsion. The first named author would like to thank Ilya Kapovich for numerous discussions on topics closely related to this paper. Many of these discussions are reflected in the present paper.
@@ 1.1 Relatively hyperbolic groups
Relatively hyperbolic groups we introduced by Gromov [23]. Various ultimately equivalent definitions have since appeared in the literature [21, 12, 26], see [27] for a comparison of the various notions. The discussion in this paper follows Osin [36].
@@
Let $G$ be a group. A finite set =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ of subgroups of $G$ is called a peripheral structure of $G$ and the elements of the $P_{i}$ are called peripheral. Any element conjugate to some peripheral element is called parabolic.
@@
Note that $X$ is only required to be a relative generating set of $G$ in Definition 1.1. In the remainder of this paper $X$ is always a generating set, in particular all relatively hyperbolic groups under consideration will be finitely generated.
@@
Osin showed that the relative Cayley graph of a relatively hyperbolic group is Gromov hyperbolic and that it fulfill the BCP-property introduced by Farb [21]. Thus Osin’s notion of relative hyperbolicity implies Farb’s notion of (strong) relative hyperbolicity.
@@ 1.2 Relatively quasiconvex subgroups
$\bar{\mathbb{O}}=\{H\cap P^{g}\mid g\in G,P\in\mathbb{P},\lvert H\cap P^{g}\rvert=\infty\}$
@@
Let $(G,\mathbb{P})$ be a relatively hyperbolic group where =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and let $X$ be a symmetric finite generating set of $G$. Let $H\leq G$ be a subgroup with finite generating set $Y\subseteq X^{*}$.
@@
It is however not the case that any structure $\mathbb{O}$ as in Theorem 1.8 for which this inclusion is a quasiisometric embedding has to be an induced structure. In fact, let $G$ be a group with finite generating set $X$ and hyperbolic relative to =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$, and $\mathbb{O}:=\{O_{1},\ldots,O_{n}\}$, where each $O_{i}$ is a finite index subgroup of $P_{i}$ and at least one $O_{i}$ is a proper subgroup of $P_{i}$. Then $G$ is not hyperbolic relative $\mathbb{O}$, since $\mathbb{O}$ is not an almost malnormal collection (cf. Osin [36], Prop. 2.36), but the inclusion $(G,d_{X\cup\mathcal{O}})\to(G,d_{X\cup\mathcal{P}})$ is a quasiisometric embedding.
@@
Let $G$ be be a group which is hyperbolic relative =$\mathbb{P}=\{P_{1},\ldots,P_{k}\}$ and $X$ a symmetric finite generating set of $G$. Let $M\geq 0$.
@@
Let $w=w_{1}\ldots w_{m}\in(Y\cup\mathcal{O})^{\ast}$ be a shortest word representing $\operatorname{lab}(s)$. Let $t$ be the path in $\operatorname{Cay}(G,X\cup\mathcal{P})$ starting in $s_{-}$ and labeled by $\iota(w)$. Since every element of $Y\cup\mathcal{O}$ is mapped by $\iota$ to a word of length at most $\mu$, it follows analogously to the proof of Theorem 1.8 that $t$ is a $(\mu\lambda,2\mu+2\mu^{2}\lambda+\mu\lambda c)$-quasigeodesic. Hence, by Lemma 1.3, $t$ has a $P_{n}$-edge $e^{\prime}$ which is connected to $e$ and whose endpoints have $X$-distance at most $\varepsilon$ from the endpoints of $e$.
@@
The claim now follows for $\nu_{H}:=\max(\bar{\nu}_{H},\varepsilon)$. $\square$
@@ 1.3 Elements of Infinite Order in Relatively Hyperbolic Groups
[…]
If $\tau(h)\geq 100\delta$, the claim follows from the minimality of $g$ and the $2\delta$-thinness of the quadrilateral $[1,g]\cup[g,hg]\cup[hg,h]\cup[h,1]$.
@@
$12d_{X\cup\mathcal{P}}(1,gP_{i})-D_{2}\leq\lvert h\rvert_{X\cup\mathcal{P}}\leq 2d_{X\cup\mathcal{P}}(1,gP_{i})+$ 1
@@
Let $(G,\mathbb{P})$ be a relatively hyperbolic group, $X$ a finite generating set of $G$. Let $\delta\geq 0$ such that $\operatorname{Cay}(G,X\cup\mathcal{P})$ is $\delta$-hyperbolic and $H$ a $\nu$-relatively quasiconvex subgroup of $G$. Let $g\in G$ such that $g$ has minimal $X\cup\mathcal{P}$-length in $gH$ and let $h\in H$. Then:
@@ 1.4 Paths with Long Peripheral Edges
Then $s$ is a $(C^{2}+3C,C^{2}+3C)$-quasigeodesic.
@@
which proves the claim. $\square$
@@ 2.1 Carrier Graphs of groups
For the basic definitions regarding graphs of groups we refer to [42] or [7], we follow the notation of [29]. Thus, the graph underlying a graph of groups $\mathbb{A}$ is $A$ and the boundary monomorphisms from an edge group $A_{e}$ into the vertex groups $A_{\alpha(e)}$ and $A_{\omega(e)}$ of the intial vertex $\alpha(e)$ and terminal vertex $\omega(e)$ of $e$ are denoted by $\alpha_{e}$ and $\omega_{e}$. An element of $\pi_{1}(\mathbb{A},v_{0})$ is represented by equivalence classes of closed $\mathbb{A}$-paths based at $v_{0}$, i.e. equivalence classes of tuples of type $(a_{0},e_{1},a_{1},\ldots,e_{k},a_{k})$ with $(e_{1},\ldots,e_{k})$ a closed path in $A$ based at $v_{0}$ and $a_{0},a_{k}\in A_{v_{0}}$ and $a_{i}\in A_{\omega}(e_{i})$ for $11\leq i\leq k-$ 1. A subgraph of groups of a graph of groups $\mathbb{A}$ is a graph of groups $\mathbb{C}$ such that $C$ is a subgraph of $A$, that $C_{x}$ is a subgroup of $A_{x}$ for all $x\in EC\cup VC$ and that the boundary monomorphism of $\mathbb{C}$ are the restrictions of the boundary monomorphisms of $\mathbb{A}$. $\mathbb{C}$ is called full if $C_{x}=A_{x}$ for all $x\in EC\cup VC$.
@@
Let $G$ be a group and $\mathcal{A}=(\mathbb{A},(g_{e})_{e\in EA},E_{0})$ be a $G$-carrier graph of groups and $v_{0}\in VA$. For any $\mathbb{A}$-path $t=(a_{0},e_{1},a_{1},\ldots,e_{k},a_{k})$ let
@@
Let $G$ be a group and $\mathcal{A}=(\mathbb{A},(g_{e})_{e\in EA},E_{0})$ a $G$-carrier graph of groups.
Let $e\in E_{0}$. A subdivision of $\mathcal{A}$ along $e$ is a $G$-carrier graph of groups $\mathcal{A}^{\prime}=(\mathbb{A}^{\prime},(g_{e})_{e\in EA^{\prime}},E^{\prime}_{0})$ with the following properties:
@@
It is obvious that such a subdivision exists for any decomposition of $g_{e}$ as described in 4. Moreover, for any $v_{0}\in VA$ there exists a canonical isomorphism $\phi\colon\pi_{1}(\mathbb{A},v_{0})\to\pi_{1}(\mathbb{A}^{\prime},v_{0})$ such that $\nu_{\mathcal{A}}=\nu_{\mathcal{A^{\prime}}}\circ\phi$.
@@
The free rank of $\mathbb{A}$ is the Betti number of the graph obtained from $A$ by collapsing all $A_{i}$ to single vertices.
@@ 2.2 $(G,\mathbb{P})$-carrier graphs of groups
Let $(G,\mathbb{P})$ be a torsion-free relatively hyperbolic group with =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and let $\mathcal{A}=(\mathbb{A},(g_{e})_{e\in EA},E_{0})$ be a $G$-carrier graph of groups. Let $C\subset A$ be a star with central vertex $c$. Let $\mathbb{C}$ be a full subgraph of groups (of $\mathbb{A}$) with underlying graph $C$ and $\mathcal{C}$ the corresponding $G$-subcarrier graph of groups.
@@
Let $G$ be a torsion-free group that is hyperbolic relative to a peripheral structure =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$. A $(G,\mathbb{P})$-carrier graph of groups is a pair
@@
    $\mathcal{A}=(\mathbb{A},(g_{e})_{e\in EA},E_{0})$ is a $G$-carrier graph of groups.
@@
be an $\mathbb{A}$-path. For any $i\leq j\in\{0,\ldots,k\}$ and $\bar{a}_{l}\in\{1,a_{l}\}$ for $l\in\{i,j\}$, the $\mathbb{A}$-path $s=(\bar{a}_{i},e_{i+1},a_{i+1},\ldots,a_{j-1},e_{j},\bar{a}_{j})$ is called an $\mathbb{A}$-subpath of $t$.
@@
    almost-simple, if $e_{i}\neq e_{j}^{\pm 1}$ for all $i\neq j\in\{1,\ldots,k\}$, and if $v_{i}=v_{j}$ for $i\neq j\in\{1,\ldots,k\}$ implies that $v_{i}$ is a central peripheral vertex.
@@
    an almost-circuit, if it is a closed almost-simple path.
@@
    $\lvert pg_{e}\rvert_{X}\geq\lvert g_{e}\rvert_{X}$ for any essential edge $e\in EA$ with $\alpha(e)\in VC_{i}$ and $p\in P_{m_{i}}$.
@@
Let now $\mathcal{A}$ be a $(G,\mathbb{P})$-carrier graph of groups such that all groups of essential vertices are quasiconvex and generated by elements of $X$-length at most $M$. Then $\mathcal{A}$ can be transformed into an $M$-prenormal carrier graph of groups by modifying the carrier graph in the way discussed below. Note that in some steps a vertex groups $A_{v}$ of some non-trivial peripheral star $\mathcal{C}$ is replaced by a strictly larger group $A^{\prime}_{v}$. Whenever this happens it is implicitly assumed that all other vertex and edge groups are also replaced by the appropriate conjugate of $A^{\prime}_{v}$ such that again all boundary monomorphisms are isomorphisms.
The (pre)normalization process is applied to each essential vertex. As the vertex group of the essential vertex is unchanged in the process the assumption on $\mathcal{A}$ guarantees that 1. and 2. of Definition 2.9 will always be satisfied.
@@
    For some maximal parabolic subgroup $U$ of $A_{v}$ there exists no essential edge $e$ such that $U$ is conjugate to $\omega_{e}(A_{e})$ in $A_{v}$.

### v2

@@ Introduction
Makanin-Razborov diagrams refine the above arguments as they give a complete parametrization of all homomorphisms from a given finitely presented (or even finitely generated) group to a given (toral relatively) hyperbolic group. They rely essentially on the ideas of Makanin [36] and Razborov [39] and are in their modern form due to Sela [42] and Kharlampovich-Myasnikov [33, 34] in the case of free groups, see [43] and [40] for hyperbolic groups and [27, 26] for toral relatively hyperbolic groups, i.e. groups that are hyperbolic relative to finitely generated free Abelian groups.
@@
Suppose that $G$ is a group that is hyperbolic relative to a collection of subgroups $\{P_{1},\ldots,P_{n}\}$. The following three settings can (and often do) yield infinitely many conjugacy classes of one-ended isomorphic subgroups of $G$:
@@
    Let $H=H_{1}*_{C}H_{2}$ be a subgroup of $G$ such that $C\leq P_{i}$ for some $i$ and let $p\in P_{i}\setminus C$ such that $p$ centralizes $C$. Then $\varphi_{p}:H\to G$ given by $\varphi_{p}(h_{1})=h_{1}$ for all $h_{1}\in H_{1}$ and $\varphi_{p}(h_{2})=ph_{2}p^{-1}$ for all $h_{2}\in H_{2}$ defines a homomorphism. It can often be ensured using a combination theorem of Martínez-Perdrosa [37] that $\varphi_{p}$ is injective and that there are infinitely many such $p$ such that images of the $\varphi_{p}$ yield isomorphic and pairwise non-conjugate subgroups isomorphic to $H$.
@@
    Let $H=H_{1}*_{C}=\langle H,t\mid tct^{-1}=\psi(c)\hbox{ for all }c\in C\rangle$ where $C\subset H_{1}\subset G$ and $C\subset P_{i}$. Let $p\in P_{i}\setminus C$ such that $p$ centralizes $C$. Then $\varphi_{p}:H\to G$ given by $\varphi_{p}(h_{1})=h_{1}$ for all $h_{1}\in H_{1}$ and $\varphi_{p}(t)=tp$ defines a homomorphism. As in (1) this construction often yields infintely many isomorphic and pairwise non-conjugate subgroups isomorphic to $H$.
@@
This is achieved by the following three steps, here $(G,\mathbb{P})$ is a torsion-free group $G$ that is hyperbolic relative to a collection of subgroups $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$.
@@
In chapter one we recall basic notions for relatively hyperbolic groups and establish a sufficient condition for a word to be a quasigeodesic. In Chapter 2 we introduce $(G,\mathbb{P})$ carrier graphs of groups and study their basic properties before we prove the combination theorem in Section 3. Section 4 is then dedicated to the introduction of folds and the proofs of the finiteness theorems. We conclude the paper with a brief discussion of the case of relatively hyperbolic groups with torsion.
The first named author would like to thank Ilya Kapovich for numerous discussions on topics closely related to this paper. Many of these discussions are reflected in the present paper. Further thanks goes to Jakob Heikamp, Nir Lazarovich and Zachary Munroe who pointed out various inaccuracies in an earlier version of this paper. In recent joint work with Edgar A. Bering IV and Jack Kohov [8] they employ the result of the present paper to study ascending chains of subgroups in relatively hyperbolic groups and 3-manifold groups.
@@ 1.1 Relatively hyperbolic groups
Relatively hyperbolic groups were introduced by Gromov [25]. Various ultimately equivalent definitions have since appeared in the literature [23, 13, 38, 28], see [29] for a comparison of the various notions. The discussion in this paper follows Osin [38].
@@
Let $G$ be a group. A finite set $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ of subgroups of $G$ is called a peripheral structure of $G$ and the elements of the $P_{i}$ are called peripheral. Any element conjugate to some peripheral element is called parabolic.
@@
Note that $X$ is only required to be a finite relative generating set of $G$ in Definition 1.1. However, in the remainder of this paper, $X$ is always a generating set, in particular all relatively hyperbolic groups under consideration will be finitely generated.
@@
Osin showed that the relative Cayley graph of a relatively hyperbolic group is Gromov hyperbolic and that it fulfills the BCP-property introduced by Farb [23]. Thus Osin’s notion of relative hyperbolicity implies Farb’s notion of (strong) relative hyperbolicity.
@@ 1.2 Relatively quasiconvex subgroups
$\bar{\mathbb{O}}=\{H\cap P^{g}\mid g\in G,P\in\mathbb{P},\lvert H\cap P^{g}\rvert$$\infty\}$
@@
Let $(G,\mathbb{P})$ be a relatively hyperbolic group where $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and let $X$ be a symmetric finite generating set of $G$. Let $H\leq G$ be a subgroup with finite generating set $Y\subseteq X^{*}$.
@@
It is however not the case that any structure $\mathbb{O}$ as in Theorem 1.8 for which this inclusion is a quasiisometric embedding has to be an induced structure. In fact, let $G$ be a group with finite generating set $X$ and hyperbolic relative to $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$, and $\mathbb{O}:=\{O_{1},\ldots,O_{n}\}$, where each $O_{i}$ is a finite index subgroup of $P_{i}$ and at least one $O_{i}$ is a proper subgroup of $P_{i}$. Then $G$ is not hyperbolic relative $\mathbb{O}$, since $\mathbb{O}$ is not an almost malnormal collection (cf. Osin [38], Prop. 2.36), but the inclusion $(G,d_{X\cup\mathcal{O}})\to(G,d_{X\cup\mathcal{P}})$ is a quasiisometric embedding.
@@
Let $G$ be be a group which is hyperbolic relative $\mathbb{P}=\{P_{1},\ldots,P_{k}\}$ and $X$ a symmetric finite generating set of $G$. Let $M\geq 0$.
@@
Let $w=w_{1}\ldots w_{m}\in(Y\cup\mathcal{O})^{\ast}$ be a shortest word representing $\operatorname{lab}(s)$. Let $t$ be the path in $\operatorname{Cay}(G,X\cup\mathcal{P})$ starting in $s_{-}$ and labeled by $\iota(w)$. Since every element of $Y\cup\mathcal{O}$ is mapped by $\iota$ to a word of length at most $\mu$. The minimality of the length of $w$ and the fact that $\mathbb{O}$ is an induced structure imply that $t$ is without backtracking. It follows analogously to the proof of Theorem 1.8 that $t$ is a $(\mu\lambda,2\mu+2\mu^{2}\lambda+\mu\lambda c)$-quasigeodesic. Hence, by Lemma 1.3, $t$ has a $P_{n}$-edge $e^{\prime}$ which is connected to $e$ and whose endpoints have $X$-distance at most $\varepsilon$ from the endpoints of $e$.
@@
The claim now follows for $\nu_{H}:=\max(\bar{\nu}_{H},2\varepsilon)$. $\square$
@@
Corollary 1.12.
Let $s=s_{1}es_{2}$ be as in Lemma 1.10 2(c), $i$ as in the conclusion of Lemma 1.10 such that $\operatorname{lab}(s_{2})\in P_{n_{i}}g_{i}$. Then there exists some $h\in H\cap P_{n_{i}}^{g_{i}}$ such that $\operatorname{lab}(s)h=\operatorname{lab}(s_{1})\hat{p}g_{i}$ with $\hat{p}\in P_{n_{i}}$ and $|\hat{p}|_{X}\leq\nu$.
Proof.
We use the notation of the proof of Lemma 1.10. Choose $p\in P_{n_{i}}$ such that $\operatorname{lab}(s_{2})=pg_{i}$. Note that $p^{\prime}:=\omega(e)\cdot\omega(e^{\prime})^{-1}\in P_{n_{i}}$ as $e$ and $e^{\prime}$ are connected. It follows that $\operatorname{lab}(s_{2}^{\prime})=\operatorname{lab}(r_{2})^{-1}(\omega(e)\cdot\omega(e^{\prime})^{-1}\operatorname{lab}(s_{2})=g_{i}^{-1}p^{\prime}pg_{i}\in H\cap P_{n_{i}}^{g_{i}}$.
The claim now follows easily with $h=\operatorname{lab}(s_{2}^{\prime})^{-1}\operatorname{lab}(\bar{s})^{-1}$ as $\operatorname{lab}(s)h=\operatorname{lab}(s_{1}^{\prime})$. $\square$
@@ 1.3 Elements of Infinite Order in Relatively Hyperbolic Groups
Since $g\in VY$, $d_{X\cup\mathcal{P}}(g,hg)\leq\tau(h)+C$ for some constant $C$ independent of $g$ and $h$. It therefore suffices to show that the difference of lengths of the (piecewise) geodesic paths $[1,h]$ and $1,g]\cup[g,hg]\cup[hg,h]$ is bounded by a constant independent of $h$.
If $\tau(h)\geq 100\delta$, the claim follows from the minimality of $g$ and the $2\delta$-thinness of the quadrilateral $1,g]\cup[g,hg]\cup[hg,h]\cup[h,1]$.
@@
$2d_{X\cup\mathcal{P}}(1,gP_{i})-D_{2}\leq\lvert h\rvert_{X\cup\mathcal{P}}\leq 2d_{X\cup\mathcal{P}}(1,gP_{i})+1$
@@
Let $(G,\mathbb{P})$ be a relatively hyperbolic group, $X$ a finite generating set of $G$. Let $\delta\geq 0$ be such that $\operatorname{Cay}(G,X\cup\mathcal{P})$ is $\delta$-hyperbolic and $H$ a $\nu$-relatively quasiconvex subgroup of $G$. Let $g\in G$ be such that $g$ has minimal $X\cup\mathcal{P}$-length in $gH$ and let $h\in H$. Then:
@@ 1.4 Paths with Long Peripheral Edges
Then $s$ is a $(C^{2}+3C,C^{2}+3C)$-quasigeodesic. Moreover $p_{i}$ and $p_{j}$ are not connected for $i\neq j\in\{1,\ldots,k\}$.
@@
which proves that $s$ is a $(C^{2}+3C,C^{2}+3C)$-quasigeodesic. That the $p_{k}$ are not connected is an immediate consequence of the fact that the $s^{N}$ is without backtracking and that $p_{k}$ is connected to $p_{k}^{N}$ for all $k$. $\square$
@@ 2.1 Carrier Graphs of groups
For the basic definitions regarding graphs of groups we refer to [44] or [7], we follow the notation of [31]. Thus, the graph underlying a graph of groups $\mathbb{A}$ is $A$ and the boundary monomorphisms from an edge group $A_{e}$ into the vertex groups $A_{\alpha(e)}$ and $A_{\omega(e)}$ of the intial vertex $\alpha(e)$ and terminal vertex $\omega(e)$ of $e$ are denoted by $\alpha_{e}$ and $\omega_{e}$. An element of $\pi_{1}(\mathbb{A},v_{0})$ is represented by equivalence classes of closed $\mathbb{A}$-paths based at $v_{0}$, i.e. equivalence classes of tuples of type $(a_{0},e_{1},a_{1},\ldots,e_{k},a_{k})$ with $(e_{1},\ldots,e_{k})$ a closed path in $A$ based at $v_{0}$ and $a_{0},a_{k}\in A_{v_{0}}$ and $a_{i}\in A_{\omega}(e_{i})$ for $1\leq i\leq k-1$. A subgraph of groups of a graph of groups $\mathbb{A}$ is a graph of groups $\mathbb{C}$ such that $C$ is a subgraph of $A$, that $C_{x}$ is a subgroup of $A_{x}$ for all $x\in EC\cup VC$ and that the boundary monomorphism of $\mathbb{C}$ are the restrictions of the boundary monomorphisms of $\mathbb{A}$. $\mathbb{C}$ is called full if $C_{x}=A_{x}$ for all $x\in EC\cup VC$.
@@
Let $G$ be a group and $\mathcal{A}=(\mathbb{A},(g_{e})_{e\in EA},E_{0},v_{0})$ be a $G$-carrier graph of groups. For any $\mathbb{A}$-path $t=(a_{0},e_{1},a_{1},\ldots,e_{k},a_{k})$ let
@@
Let $G$ be a group and $\mathcal{A}=(\mathbb{A},(g_{e})_{e\in EA},E_{0},v_{0})$ a $G$-carrier graph of groups.
Let $e\in E_{0}$. A subdivision of $\mathcal{A}$ along $e$ is a $G$-carrier graph of groups $\mathcal{A}^{\prime}=(\mathbb{A}^{\prime},(g_{e})_{e\in EA^{\prime}},E^{\prime}_{0},v_{0}^{\prime})$ with the following properties:
@@
  7. 7.
    $v_{0}^{\prime}=v_{0}$.
It is obvious that such a subdivision exists for any decomposition of $g_{e}$ as described in 4. Moreover there exists a canonical isomorphism $\phi\colon\pi_{1}(\mathbb{A},v_{0})\to\pi_{1}(\mathbb{A}^{\prime},v_{0}^{\prime})$ such that $\nu_{\mathcal{A}}=\nu_{\mathcal{A^{\prime}}}\circ\phi$.
@@
The relative rank of $\mathbb{A}$ is the Betti number of the graph obtained from $A$ by collapsing all $A_{i}$ to single vertices.
@@ 2.2 $(G,\mathbb{P})$-carrier graphs of groups
Let $(G,\mathbb{P})$ be a torsion-free relatively hyperbolic group with $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and let $\mathcal{A}=(\mathbb{A},(g_{e})_{e\in EA},E_{0})$ be a $G$-carrier graph of groups. Let $C\subset A$ be a star with central vertex $c$. Let $\mathbb{C}$ be a full subgraph of groups (of $\mathbb{A}$) with underlying graph $C$ and $\mathcal{C}$ the corresponding $G$-subcarrier graph of groups.
@@
Let $G$ be a torsion-free group that is hyperbolic relative to a peripheral structure $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$. A $(G,\mathbb{P})$-carrier graph of groups is a pair
@@
    $\mathcal{A}=(\mathbb{A},(g_{e})_{e\in EA},E_{0},v_{0})$ is a finite $G$-carrier graph of groups.
@@
  7. 7.
    $\mathbb{A}$ is minimal relative to base point and relative to the non-trivial central peripheral vertices, i.e. for any valence $1$ vertex $v\in VA$ and $e\in EA$ with $\alpha(e)=v$ the boundary morphism $\alpha_{e}$ is not surjective unless $v$ is a non-trivial central peripheral vertex or the base vertex.
@@
be an $\mathbb{A}$-path. For any $i\leq j\in\{0,\ldots,k\}$ and $\bar{a}_{l}\in\{1,a_{l}\}$ for $l\in\{i,j\}$, the $\mathbb{A}$-path $s=(\bar{a}_{i},e_{i+1},a_{i+1},\ldots,a_{j-1},e_{j},\bar{a}_{j})$ is called an $\mathbb{A}$-subpath of $t$. $s$ is called a full $\mathbb{A}$-subpath if $\bar{a}_{l}=a_{l}$ for $l\in\{i,f\}$.
@@
    $p$ is called almost-simple, if the following hold:
    1. (a)
      $e_{i}\neq e_{j}^{\pm 1}$ for all $i\neq j\in\{1,\ldots,k\}$.
    2. (b)
      $v_{i}=v_{j}$ for $i\neq j\in\{0,\ldots,k\}$ implies that $v_{i}$ is a central peripheral vertex.
@@
    $p$ is called an almost-circuit, if $p$ is closed and the following hold:
    1. (a)
      $e_{i}\neq e_{j}^{\pm 1}$ for all $i\neq j\in\{1,\ldots,k\}$.
    2. (b)
      $v_{i}=v_{j}$ for $i\neq j\in\{0,\ldots,k\}$ implies that $\{i,j\}=\{0,m\}$ or that $v_{i}$ is a central peripheral vertex.
@@
    $\lvert pg_{e}a\rvert_{X}\geq\lvert g_{e}\rvert_{X}$ for any essential edge $e\in EA$ with $\alpha(e)\in VC_{i}$, $p\in P_{m_{i}}$ and $a\in A_{v}$.
@@
Let now $\mathcal{A}=(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ be a $(G,\mathbb{P})$-carrier graph of groups (we will continue to use this shorthand in the context of $(G,\mathbb{P})$-carrier graphs of groups) such that all groups of essential vertices are quasiconvex and generated by elements of $X$-length at most $M$. Then $\mathcal{A}$ can be transformed into an $M$-prenormal carrier graph of groups by modifying the carrier graph in the way discussed below. In eacht step of this construction the following is assumed:
  1. 1.
    If a vertex group $A_{v}$ of some non-trivial peripheral star $\mathcal{C}$ is replaced by a strictly larger group $A^{\prime}_{v}$ then all other vertex and edge groups of that peripheral star are also replaced by the appropriate conjugate of $A^{\prime}_{v}$ such that again all boundary monomorphisms are isomorphisms.
  2. 2.
    The orientation $E_{0}$ induces a natural orientation of the new $\mathbb{A}$-graph with possibly emerging edges being oriented towards essential vertices.
  3. 3.
    The image of the base vertex is the base vertex of the new $\mathbb{A}$-graph.
The (pre)normalization process is applied to each essential vertex. It follows from results of the second author [17], that this process can be carried out effectively in many relevant situations. As the vertex group of the essential vertex is unchanged in the process the assumption on $\mathcal{A}$ guarantees that 1. and 2. of Definition 2.9 will always be satisfied.
@@
    For some maximal parabolic subgroup $U$ of $A_{v}$ there exists no essential edge $e$ such that $\omega_{e}(A_{e})$ is in $A_{v}$ conjugate to a subgroup of $U$.